OpenStudy (anonymous):

THE LENGTH OF A RECTANGLE IS 6 TIMES THE WIDTH. THE PERIMETER OF THE RECTANGLE IS 105. WHAT IS THE LENGTH AND WIDTH OF THE RECTANGLE?

OpenStudy (anonymous):

don.t understand

OpenStudy (anonymous):

L = 6w so P = the sum of the sides , then we have : P = 6w + 6w + w + w 105 = 14w so w = 15/2 and L = 2(w) = 2(15/2) = 15 ^_^ I guess

OpenStudy (anonymous):

L=6W= 6(15/2)=45

OpenStudy (anonymous):

alina, this answer is fine, except sstarica just made a slight error at the end which AnwarA corrected for. sstarica had everything right up to the line, w=15/2. After that, where he/she has written L=2(w), that's wrong, but it's a typo since he/she had the relationship between L and w right at the beginning. The width calculated by sstarica is correct. Then take the width and use L=6w like AnwarA did. Is that okay?

OpenStudy (anonymous):

oops! my bad >_< thanks loki + anwar